A circular race course has an inner diameter of 318 m and an outer diameter of 324 m. What is the average length of the course to the nearest meter?
Answer by palanisamy(496)

Given,A circular race course has an inner diameter of 318 m and an outer diameter of 324 m.
Therefore the average diameter = (318+324)/2 = 321 m
The average length of the course = (22/7)*321 = 1009 m nearly